OpenStudy (anonymous):

Write the slope of a line that is perpendicular to each line: x+7y=-21

hero (hero):

Saifoo's answer to that previous one was wrong

OpenStudy (anonymous):

what was the answer?

hero (hero):

Of course, you deleted it before I could point out the discrepancy

OpenStudy (anonymous):

solve for y \[y = -\frac{1}{7}x - 3\] where m = -1/7 in order for a slope to be perpendicular to another they must add up to -1 so \[-\frac{1}{7} + x = -1\] so the new slope would be 7

OpenStudy (anonymous):

not when they are added.. but when they are multiplied.. so it should be \[-\frac{1}{7}x = -1\]